OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (e.mccormick):

What have you tried to get this?

OpenStudy (e.mccormick):

If you find the slope, you can use the slope and one point to define the line.

OpenStudy (texaschic101):

m = (y2 - y1)/ (x2 - x1) -- m is the slope...set 1(-2,3)..set 2 (2,7) m = (7 - 3) / (2 - (- 2) m = 4/4 = 1 slope is 1 y - y1 = m(x - x1) ....I am using points (-2,3) y - 3 = 1(x - (-2) y - 3 = 1x + 2 y = 1x + 2 + 3 y = x + 5 x - y = -5